OpenStudy (emmily_cool):

The rectangular prism shown has a volume of 140 m3. This prism is enlarged using a scale factor of 2. What is the volume of the enlarged prism? A. 280 m3 B. 560 m3 C. 1120 m3 D. 1260 m3

OpenStudy (emmily_cool):

@threshsupport

OpenStudy (mjdennis):

how do you find the volume of a rectangular prism? v= L*w*h, right?

OpenStudy (emmily_cool):

ya

OpenStudy (mjdennis):

so L*w*h = 140, but we don't know L*w*h let's scale everything up by 2: V' = 2*L * 2*w * 2*h And group like this: V' = 2*2*2 * L*w*h Good so far?

OpenStudy (emmily_cool):

ya

OpenStudy (mjdennis):

So, substitute in a number in place of L*w*h. What did we just decide L*w*h is? V' = 8*______ = your answer!

OpenStudy (mjdennis):

@codyhelton75 , go read the Code of Conduct. Not only are you violating it, you can't prove your answer is right. http://openstudy.com/code-of-conduct

OpenStudy (codyhelton75):

it said scale by to so 140x2=280

OpenStudy (mjdennis):

And scale by 2 usually means each dimension goes up by 2. But I could be wrong here. Still, if you are contributing to a solution, your job is to help the asker find the answer, not give it away. Do you want to talk to a moderator about it, instead of me?